OpenStudy (soggylearnsandhelps):

Write the equation in standard form. Then factor the left side of the equation. 2x^2 + 7x = 15

OpenStudy (3mar):

May I help?

OpenStudy (soggylearnsandhelps):

Yeah! My answer is (2x + 3)(x + 5) = 0

OpenStudy (soggylearnsandhelps):

I wanna see if it correct or not,

OpenStudy (3mar):

what is the middle teem after you factored?

OpenStudy (soggylearnsandhelps):

Eh?

OpenStudy (3mar):

the coefficient of the term x?

OpenStudy (soggylearnsandhelps):

Hold on let me write my solution down,

OpenStudy (soggylearnsandhelps):

or the way i solved it,

OpenStudy (3mar):

It'd better!

OpenStudy (soggylearnsandhelps):

x^2+7x-15=0 Factor 2x^2+7x-15:(x+5)(2x-3)

OpenStudy (soggylearnsandhelps):

(x+5)(2x-3)=0

OpenStudy (soggylearnsandhelps):

Wait thats my answer right there,

OpenStudy (soggylearnsandhelps):

(2x + 5)(x – 3) = 0 (2x + 3)(x + 5) = 0 (2x – 3)(x + 5) = 0 (2x – 5)(x + 3) = 0

OpenStudy (soggylearnsandhelps):

those are the answers,

OpenStudy (soggylearnsandhelps):

My answer is C

OpenStudy (3mar):

Great! That makes sense now!

OpenStudy (3mar):

Because you mentioned earlier "Yeah! My answer is (2x + 3)(x + 5) = 0"!! and this is not correct!

OpenStudy (soggylearnsandhelps):

Yeah I had to redo it in my head, XD

OpenStudy (soggylearnsandhelps):

Ill give ya a medal XD just for fun!

OpenStudy (3mar):

Congratulations!

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (soggylearnsandhelps):

Thanks for the help,

OpenStudy (3mar):

You know!, Medals don't matter. What matters is that you get the idea and can do solving such this yourself.

OpenStudy (soggylearnsandhelps):

Yeah, I did solve it myself but I just wanted to reward you because you came on and helped a little!

OpenStudy (soggylearnsandhelps):

Thanks!

OpenStudy (3mar):

Thank you for learning!!

OpenStudy (3mar):

No problem at all!

OpenStudy (3mar):

Any more questions?

OpenStudy (soggylearnsandhelps):

Um, No. If i do ill post them on here with my answer so it can be checked,

OpenStudy (soggylearnsandhelps):

Thanks again!

OpenStudy (3mar):

Any Help... Anytime. You are welcome!